# Quantum conditional entropies and fully entangled fraction of states with maximally mixed marginals

### v1

@@ III.2 Weyl state
The conditional Rényi 2-entropy of $\rho^{weyl}_{2}$ is given by $S_{2}(A|B)_{\rho^{weyl}_{2}}=1-\log_{2}(1+t_{1}^{2}+t_{2}^{2}+t_{3}^{2})$ and the fully entangled fraction of Weyl state is given by $FEF(\rho^{weyl}_{2})=\frac{1}{4}[1+N(\rho^{weyl}_{2})]$, where $N(\rho^{weyl}_{2})=Tr|T|$ and $T$ is correlation matrix.
@@
$\Rightarrow FEF(\rho^{weyl}_{2})>\frac{1}{2}{}$
@@
Proof: Let FEF(ρ2weyl) >$\text{{\bf{Proof:}}\leavevmode\nobreak\ Let $FEF(\rho^{weyl}_{2})$ $>\frac{1}{2}$}\Rightarrow N(\rho^{weyl}_{2})>1{}$
@@
$\Rightarrow S_{2}(A|B)_{\rho^{weyl}_{2}}<\log_{2}\left[\frac{1}{1-R}\right]{}$
@@ IV.1 Isotropic states
where $|\psi_{d}^{+}\rangle=\frac{1}{\sqrt{d}}\sum_{i=1}|ii\rangle$ and $FEF(\rho_{d}^{iso})=F$. Conditional von Neumann entropy (CVNE) is given by
$S(A|B)_{\rho^{iso}_{d}}=$ $-F\log_{2}(F)-(1-F)\log_{2}\left(\frac{1-F}{d^{2}-1}\right){}$ (19)
$-\log_{2}(d).$
@@ IV.1.1 Relation with conditional von Neumann entropy
$\Rightarrow F>\log_{(d^{2}-1)}\left[\frac{d^{2}-1}{d}\right]{}$
@@
If conditional von Neumann entropy (CVNE) of isotropic state is negative then fully entangled fraction of isotropic state is strictly greater than $\frac{1}{d}$ i.e.,CVNE$(\rho_{d}^{iso})<0\Rightarrow F>\frac{1}{d}$
@@ IV.1.2 Relation with conditional Rényi 2-entropy
$\Leftrightarrow F>\frac{1+\sqrt{1+d(d^{2}-d-1)}}{d^{2}}{}$
@@ IV.2 Werner state in two qudits
where $V$ is swap operator such that $V(|ij\rangle)=|ji\rangle.$ FEF for Werner states is given in [35].
@@ IV.2.1 Relation with conditional von Neumann entropy
$S(A|B)=$ $-\frac{1+x}{2}\log_{2}\frac{1+x}{d^{2}+d}{}$ (22)
$-\frac{1-x}{2}\log_{2}\frac{1-x}{d^{2}-d}-\log_{2}d.$
@@ IV.2.2 Relation with conditional Rényi $\alpha$-entropy
$S_{\alpha}(A|B)=$ $\frac{1}{1-\alpha}\log_{2}[d^{\alpha-1}[\left(\frac{d^{2}+d}{2}\right)\left(\frac{1+x}{d^{2}+d}\right)^{\alpha}{}$ (23)
$+\left(\frac{d^{2}-d}{2}\right)\left(\frac{1-x}{d^{2}-d}\right)^{\alpha}]],\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \alpha>1$
@@
$S_{\alpha}(A|B)=\frac{1}{1-\alpha}\log_{2}[d^{\alpha-1}[\left(\frac{d^{2}+d}{2}\right)\left(\frac{1+x}{d^{2}+d}\right)^{\alpha}{}$
$+\left(\frac{d^{2}-d}{2}\right)\left(\frac{1-x}{d^{2}-d}\right)^{\alpha}]]{}$
$\Rightarrow 2^{(1-\alpha)S_{\alpha}(A|B)}={}$
$\frac{d^{\alpha-1}(d^{2}+d)F^{\alpha}+d^{\alpha-1}(d^{2}-d)^{1-\alpha}(1-x)^{\alpha}}{2}{}$
$\Rightarrow F^{\alpha}=\frac{2.2^{(1-\alpha)S_{\alpha}(A|B)}-d^{\alpha-1}(d^{2}-d)^{1-\alpha}(1-x)^{\alpha}}{d^{\alpha-1}(d^{2}+d)}{}$
@@
If $d$ is even or odd and $\frac{1}{d}\leq x\leq 1$, then $F^{\alpha}=\Delta$ where $\Delta=\frac{2.2^{(1-\alpha)S_{\alpha}(A|B)}-d^{\alpha-1}(d^{2}+d)^{1-\alpha}(1+x)^{\alpha}}{d^{\alpha-1}(d^{2}-d)}$.
@@
$S_{\alpha}(A|B)=\frac{1}{1-\alpha}\log_{2}[d^{\alpha-1}[\left(\frac{d^{2}+d}{2}\right)\left(\frac{1+x}{d^{2}+d}\right)^{\alpha}{}$
$+\left(\frac{d^{2}-d}{2}\right)\left(\frac{1-x}{d^{2}-d}\right)^{\alpha}]]{}$
$\Rightarrow 2^{(1-\alpha)S_{\alpha}(A|B)}={}$
$\frac{d^{\alpha-1}(d^{2}+d)(1+x)^{\alpha}+d^{\alpha-1}(d^{2}-d)^{1-\alpha}F^{\alpha}}{2}{}$
$\Rightarrow F^{\alpha}=\frac{2.2^{(1-\alpha)S_{\alpha}(A|B)}-d^{\alpha-1}(d^{2}+d)^{1-\alpha}(1+x)^{\alpha}}{d^{\alpha-1}(d^{2}-d)}{}$
@@ IV.3 Rank-deficient states
$\rho_{rd}=p|\psi_{d}^{+}\rangle\langle\psi_{d}^{+}|+(1-p)|01\rangle\langle 01|,p\in(0,1]$ (24)
@@ IV.3.1 Relation with conditional von Neumann entropy
$S(A|B)_{\rho_{rd}}=\frac{dp-p}{d}\log_{2}\left(\frac{p}{d}\right){}$
$+\frac{d-dp+p}{d}\log_{2}\left(\frac{d-dp+p}{d}\right)+H(p),$ (25)
@@ IV.4.1 Relation with conditional von Neumann entropy
If $F>\frac{1}{d}$ then $S(A|B)<Y$ where $Y=\log_{2}\left(\frac{d^{F-1}}{\prod_{i=0}^{d^{2}-2}p_{i}^{p_{i}}}\right),p_{i}>0$.
@@ IV.4.2 Relation with other $\alpha-$ conditional entropies
Conditional Rényi $\alpha$- entropy (CRAE) of $\rho_{gb}$ is given by $S_{\alpha}(A|B)=\frac{1}{1-\alpha}\log_{2}\left(d^{\alpha-1}\sum_{i=0}^{d^{2}-1}p_{i}^{\alpha}\right)\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \alpha>1$
@@
$\Rightarrow F^{\alpha}+\sum_{i=0}^{d^{2}-2}p_{i}^{\alpha}>\frac{1}{d^{\alpha-1}}\Rightarrow F^{\alpha}>\frac{1}{d^{\alpha-1}}-\sum_{i=0}^{d^{2}-2}p_{i}^{\alpha}$
@@
If $F>\frac{1}{d}$ then $S_{\alpha}(A|B)<\frac{1}{\alpha-1}\log_{2}\left(\frac{1+d^{\alpha}Z}{d}\right)$ where $Z=\sum_{i=0}^{d^{2}-2}p_{i}^{\alpha}$ i.e., sum of all $p_{i}^{\alpha}$ except the term for largest $p_{i}$.
@@
$\Rightarrow S_{\alpha}(A|B)<\frac{1}{1-\alpha}\log_{2}\left(\frac{1+d^{\alpha}Z}{d}\right){}$
@@ VI.1 Weyl states
By using Theorem (4), we know that two-qubit Weyl states with negative CR2222E are useful for teleportation i.e., $FEF(\rho^{weyl}_{2})>\frac{1}{2}$, and from [39] we know that all quantum states useful for teleportation are nonlocal resources and hence $k-$copy nonlocal. Thus, we have the following result.
@@ VI.3 Noisy state
We consider an arbitrary pure state $|\psi\rangle$ with its Schmidt decomposition given by
@@
In the below, we see the relation between $k-$copy nonlocal and negative conditional entropies for $d=3,4,5$. Here, we are fixing the Schmidt coefficient of state $\delta_{d}^{ns}$ in Equation (31) for respective dimension. Considering the Schmidt coefficient ($\lambda_{0}=\frac{1}{\sqrt{2}},\lambda_{1}=\frac{1}{\sqrt{3}},\lambda_{2}=\frac{1}{\sqrt{6}}$), ($\lambda_{0}=\lambda_{1}=\frac{1}{\sqrt{3}},\lambda_{2}=\lambda_{3}=\frac{1}{\sqrt{6}}$) and ($\lambda_{0}=\frac{1}{\sqrt{3}},\lambda_{1}=\frac{1}{\sqrt{7}},\lambda_{2}=\sqrt{\frac{2}{21}},\lambda_{3}=\sqrt{\frac{4}{21}},\lambda_{4}=\sqrt{\frac{5}{21}}$) for dimension $d=3,4$ and $5$, respectively. We observe that the state is $k-$copy nonlocal and have negative CR2222E in the range ($p>0.263305$ $p>0.516398$), ($p>0.206292$ $p>0.45399$) and ($p>0.174342$ $p>0.415577$) for dimension $d=3,4$ and $5$, respectively. Thus, we can say that if the state $\delta_{d}^{ns}(d=3,4,5)$ has negative CR2222E for the above restriction, then it is $k-$copy nonlocal. Also, we observe that the state is $k-$copy nonlocal and have negative CVNE in the range ($p>0.263305$ $p>0.728901$), ($p>0.206292$ $p>0.699086$) and ($p>0.174342$ $p>0.685898$) for dimension $d=3,4$ and $5$, respectively. Thus, we can say that if the state $\delta_{d}^{ns}(d=3,4,5)$ has negative CVNE for the above restriction, then it is $k-$copy nonlocal
VII Application: Work Extraction
Consider any quantum state $\rho\in\mathfrak{B}(\mathbb{H}_{AB})$ where $\mathbb{H}_{AB}=\mathbb{H}_{A}\otimes\mathbb{H}_{B}$, $\text{dim}\leavevmode\nobreak\ \mathbb{H}_{A}=\text{dim}\leavevmode\nobreak\ \mathbb{H}_{B}=d$. In [68], the author introduced a framework and defined minimal deterministic average work cost, denoted by $W_{g}(\rho)$, in the IID (independent and identically distributed) limit of the erasure process according to [51]. Upper bounds for $W_{g}(\rho)$ is given by [51, 68]
@@ VII Application: Work Extraction
VII.0.1 Werner state in two qubits
@@
where $\delta=\text{min}\{\delta_{1},\delta_{2}\}$, $\delta_{i}\text{ are eigenvalues of }\rho^{wer}_{2}$ and the multiplicity of eigenvalue $\delta_{2}$ is three.
VII.0.2 Generalized Bell Diagonal States

### v2

@@ III.2 Weyl state
The conditional Rényi 2-entropy of $\rho^{weyl}_{2}$ is given by $S_{2}(A|B)_{\rho^{weyl}_{2}}=1-\log_{2}(1+t_{1}^{2}+t_{2}^{2}+t_{3}^{2})$ and the fully entangled fraction of Weyl state is given by $\text{FEF}(\rho^{weyl}_{2})=\frac{1}{4}[1+N(\rho^{weyl}_{2})]$, where $N(\rho^{weyl}_{2})=Tr|T|$ and $T$ is correlation matrix.
@@
$\Rightarrow\text{FEF}(\rho^{weyl}_{2})>\frac{1}{2}.{}$
@@
Proof: Let FEF(ρ2weyl) >$\text{{\bf{Proof:}}\leavevmode\nobreak\ Let $\text{FEF}(\rho^{weyl}_{2})$ $>\frac{1}{2}$}\Rightarrow N(\rho^{weyl}_{2})>1{}$
@@
$\Rightarrow S_{2}(A|B)_{\rho^{weyl}_{2}}<\log_{2}\left[\frac{1}{1-R}\right].{}$
@@ IV.1 Isotropic states
where $|\psi_{d}^{+}\rangle=\frac{1}{\sqrt{d}}\sum_{i=1}|ii\rangle$ and $\text{FEF}(\rho_{d}^{iso})=F$. Conditional von Neumann entropy (CVNE) is given by
$S(A|B)_{\rho^{iso}_{d}}=-F\log_{2}(F)-(1-F)\log_{2}\left(\frac{1-F}{d^{2}-1}\right)-\log_{2}(d).$ (19)
@@ IV.1.1 Relation with conditional von Neumann entropy
$\Rightarrow F>\log_{(d^{2}-1)}\left[\frac{d^{2}-1}{d}\right].{}$
@@
If conditional von Neumann entropy (CVNE) of isotropic state is negative then fully entangled fraction of isotropic state is strictly greater than $\frac{1}{d}$ i.e., CVNE$(\rho_{d}^{iso})<0\Rightarrow F>\frac{1}{d}.$
@@ IV.1.2 Relation with conditional Rényi 2-entropy
$\Leftrightarrow F>\frac{1+\sqrt{1+d(d^{2}-d-1)}}{d^{2}}.{}$
@@ IV.2 Werner state in two qudits
where $V$ is swap operator such that $V(|ij\rangle)=|ji\rangle.$ FEF for Werner states is given in [35], and denoted by $F$.
@@ IV.2.1 Relation with conditional von Neumann entropy
$S(A|B)=-\frac{1+x}{2}\log_{2}\frac{1+x}{d^{2}+d}-\frac{1-x}{2}\log_{2}\frac{1-x}{d^{2}-d}-\log_{2}d.$ (22)
@@ IV.2.2 Relation with conditional Rényi $\alpha$-entropy
$S_{\alpha}(A|B)=$ $\frac{1}{1-\alpha}\log_{2}[d^{\alpha-1}[\left(\frac{d^{2}+d}{2}\right)\left(\frac{1+x}{d^{2}+d}\right)^{\alpha}+\left(\frac{d^{2}-d}{2}\right)\left(\frac{1-x}{d^{2}-d}\right)^{\alpha}]],\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \alpha>1.$ (23)
@@
$S_{\alpha}(A|B)=\frac{1}{1-\alpha}\log_{2}[d^{\alpha-1}[\left(\frac{d^{2}+d}{2}\right)\left(\frac{1+x}{d^{2}+d}\right)^{\alpha}+\left(\frac{d^{2}-d}{2}\right)\left(\frac{1-x}{d^{2}-d}\right)^{\alpha}]]{}$
$\Rightarrow 2^{(1-\alpha)S_{\alpha}(A|B)}=\frac{d^{\alpha-1}(d^{2}+d)F^{\alpha}+d^{\alpha-1}(d^{2}-d)^{1-\alpha}(1-x)^{\alpha}}{2}{}$
$\Rightarrow F^{\alpha}=\frac{2.2^{(1-\alpha)S_{\alpha}(A|B)}-d^{\alpha-1}(d^{2}-d)^{1-\alpha}(1-x)^{\alpha}}{d^{\alpha-1}(d^{2}+d)}.{}$
@@
If $d$ is even and $-1\leq x<\frac{1}{d}$, then $F^{\alpha}=\Delta$ where $\Delta=\frac{2.2^{(1-\alpha)S_{\alpha}(A|B)}-d^{\alpha-1}(d^{2}+d)^{1-\alpha}(1+x)^{\alpha}}{d^{\alpha-1}(d^{2}-d)}$.
@@
$S_{\alpha}(A|B)=\frac{1}{1-\alpha}\log_{2}[d^{\alpha-1}[\left(\frac{d^{2}+d}{2}\right)\left(\frac{1+x}{d^{2}+d}\right)^{\alpha}+\left(\frac{d^{2}-d}{2}\right)\left(\frac{1-x}{d^{2}-d}\right)^{\alpha}]]{}$
$\Rightarrow 2^{(1-\alpha)S_{\alpha}(A|B)}=\frac{d^{\alpha-1}(d^{2}+d)^{1-\alpha}(1+x)^{\alpha}+d^{\alpha-1}(d^{2}-d)F^{\alpha}}{2}{}$
$\Rightarrow F^{\alpha}=\frac{2.2^{(1-\alpha)S_{\alpha}(A|B)}-d^{\alpha-1}(d^{2}+d)^{1-\alpha}(1+x)^{\alpha}}{d^{\alpha-1}(d^{2}-d)}.{}$
@@ IV.3 Rank-deficient states
$\rho_{rd}=p|\psi_{d}^{+}\rangle\langle\psi_{d}^{+}|+(1-p)|01\rangle\langle 01|,\leavevmode\nobreak\ \leavevmode\nobreak\ p\in(0,1].$ (24)
@@ IV.3.1 Relation with conditional von Neumann entropy
$S(A|B)_{\rho_{rd}}=\frac{dp-p}{d}\log_{2}\left(\frac{p}{d}\right)+\frac{d-dp+p}{d}\log_{2}\left(\frac{d-dp+p}{d}\right)+H(p),$ (25)
@@ IV.4.1 Relation with conditional von Neumann entropy
If $F>\frac{1}{d}$ then $S(A|B)<Y$ where $Y=\log_{2}\left(\frac{d^{F-1}}{\prod_{i=0}^{d^{2}-2}p_{i}^{p_{i}}}\right)$, and $\prod_{i=0}^{d^{2}-2}p_{i}^{p_{i}}$ is product of all $p_{i}^{p_{i}}$ except the term for largest $p_{i}$, $p_{i}>0$.
@@ IV.4.2 Relation with other $\alpha-$ conditional entropies
Conditional Rényi $\alpha$- entropy (CRAE) of $\rho_{gb}$ is given by $S_{\alpha}(A|B)=\frac{1}{1-\alpha}\log_{2}\left(d^{\alpha-1}\sum_{i=0}^{d^{2}-1}p_{i}^{\alpha}\right)\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \alpha>1.$
@@
$\Rightarrow F^{\alpha}+\sum_{i=0}^{d^{2}-2}p_{i}^{\alpha}>\frac{1}{d^{\alpha-1}}\Rightarrow F^{\alpha}>\frac{1}{d^{\alpha-1}}-\sum_{i=0}^{d^{2}-2}p_{i}^{\alpha}.$
@@
If $F>\frac{1}{d}$ then $S_{\alpha}(A|B)<\frac{1}{1-\alpha}\log_{2}\left(\frac{1+d^{\alpha}Z}{d}\right)$ where $Z=\sum_{i=0}^{d^{2}-2}p_{i}^{\alpha}$ i.e., sum of all $p_{i}^{\alpha}$ except the term for largest $p_{i}$.
@@
$\Rightarrow S_{\alpha}(A|B)<\frac{1}{1-\alpha}\log_{2}\left(\frac{1+d^{\alpha}Z}{d}\right).{}$
@@ VI.1 Weyl states
By using Theorem (4), we know that two-qubit Weyl states with negative CR2222E are useful for teleportation i.e., $\text{FEF}(\rho^{weyl}_{2})>\frac{1}{2}$, and from [39] we know that all quantum states useful for teleportation are nonlocal resources and hence $k-$copy nonlocal. Thus, we have the following result.
@@ VI.3 Noisy state
We consider a pure state $|\psi\rangle$ with its Schmidt decomposition given by
@@
In the below, we see the relation between $k-$copy nonlocal and negative conditional entropies for $d=3,4,5$. Here, we are fixing the Schmidt coefficient of state $\delta_{d}^{ns}$ in Equation (31) for respective dimension. Considering the Schmidt coefficient ($\lambda_{0}=\frac{1}{\sqrt{2}},\lambda_{1}=\frac{1}{\sqrt{3}},\lambda_{2}=\frac{1}{\sqrt{6}}$), ($\lambda_{0}=\lambda_{1}=\frac{1}{\sqrt{3}},\lambda_{2}=\lambda_{3}=\frac{1}{\sqrt{6}}$) and ($\lambda_{0}=\frac{1}{\sqrt{3}},\lambda_{1}=\frac{1}{\sqrt{7}},\lambda_{2}=\sqrt{\frac{2}{21}},\lambda_{3}=\sqrt{\frac{4}{21}},\lambda_{4}=\sqrt{\frac{5}{21}}$ ) for dimension $d=3,4$ and $5$, respectively. We observe that the state is $k-$copy nonlocal and have negative CR2222E in the range ($p>0.263305$ $p>0.516398$), ($p>0.206292$ $p>0.45399$) and ($p>0.174342$ $p>0.415577$) for dimension $d=3,4$ and $5$, respectively. Thus, we can say that if the state $\delta_{d}^{ns}(d=3,4,5)$ has negative CR2222E for the above restriction, then it is $k-$copy nonlocal. Also, we observe that the state is $k-$copy nonlocal and have negative CVNE in the range ($p>0.263305$ $p>0.728901$), ($p>0.206292$ $p>0.699086$) and ($p>0.174342$ $p>0.685898$) for dimension $d=3,4$ and $5$, respectively. Thus, we can say that if the state $\delta_{d}^{ns}(d=3,4,5)$ has negative CVNE for the above restriction, then it is $k-$copy nonlocal.
VII Application
Having seen some implications of the relations in the context of multi-copy nonlocality, we now proceed to provide some applications in this section.
VII.1 Work Extraction
Consider any quantum state $\rho\in\mathfrak{B}(\mathbb{H}_{AB})$ where $\mathbb{H}_{AB}=\mathbb{H}_{A}\otimes\mathbb{H}_{B}$, $\text{dim}\leavevmode\nobreak\ \mathbb{H}_{A}=\text{dim}\leavevmode\nobreak\ \mathbb{H}_{B}=d$. In [75], the authors introduced a framework and defined minimal deterministic average work cost, denoted by $W_{g}(\rho)$, in the IID (independent and identically distributed) limit of the erasure process according to [51]. The authors in [75] used the twirling operation and FEF of isotropic states to achieve bounds to $W_{g}(\rho)$ in terms of FEF. We obtain bounds for $W_{g}(\rho)$, using the relations obtained between FEF and QCE obtained in our work, for some special class of states detailed below.
Bounds for $W_{g}(\rho)$ is given by [51, 75]
@@ VII.1 Work Extraction
VII.1.1 Werner state in two qubits
@@
where $\delta=\text{min}\{\delta_{1},\delta_{2}\}$, $\delta_{1}=\frac{1-p}{4}$  and  $\delta_{2}=\frac{1+3p}{4}$ are eigenvalues of $\rho^{wer}_{2}$ with multiplicity $3$ and $1$, respectively.
VII.1.2 Generalized Bell Diagonal States
@@
VII.2 Faithful entanglement
Detection of entanglement in quantum systems is a core problem in quantum information science. Negativity of conditional entropy is a signature of entanglement. Within the class of entangled states there are certain states which can be detected by a proximity-measure based witness. Such proximity-measured witnesses are usually constructed from FEF. In this subsection, we show that such entangled states can also be detected through conditional entropies, which is a consequence of the relations obtained between FEF and CE.
Consider any quantum state $\rho\in\mathfrak{B}(\mathbb{H}_{AB})$ where $\mathbb{H}_{AB}=\mathbb{H}_{A}\otimes\mathbb{H}_{B}$, $\text{dim}\leavevmode\nobreak\ \mathbb{H}_{A}=\text{dim}\leavevmode\nobreak\ \mathbb{H}_{B}=d$. Now, consider a pure entangled state $|\psi\rangle$ such that the fidelity of $\rho$ with respect to pure entangled state $|\psi\rangle$ is denoted by $F_{|\psi\rangle}(\rho)$, and is given as $F_{|\psi\rangle}(\rho)=\langle\psi|\rho|\psi\rangle$. Fidelity is a measure of proximity of a state $\rho$ with a pure entangled state $|\psi\rangle$. The fidelity-based entanglement witness can be given as $W_{|\psi\rangle}=\eta I-|\psi\rangle\langle\psi|$ for some suitably chosen real number $\eta$. On measuring the observable $W_{|\psi\rangle}$, the quantity $\textbf{Tr}(\rho W_{|\psi\rangle})$ is obtained. If $F_{|\psi\rangle}(\rho)$ is greater than some threshold value $\eta$, then the witness operator detects the presence of entanglement in $\rho$. Such entangled states are termed as faithful [52, 53, 54]. If the state $|\psi\rangle$ is a maximally entangled pure state i.e., $|\psi\rangle=|\phi^{+}\rangle=\frac{1}{\sqrt{d}}\sum_{i=1}^{d}|ii\rangle$, then the smallest possible value of $\eta$ is $\frac{1}{d}$, and the witness operator is given as $W_{|\psi\rangle}=\frac{I}{d}-|\psi\rangle\langle\psi|$. From [54], we know that a state $\rho$ is faithful if and only if $\text{FEF}(\rho)>\frac{1}{d}$. Since, faithful entanglement is directly related to FEF, we can now characterize faithful entanglement in terms of conditional entropies. Since, direct evaluation of FEF is computationally difficult albeit for some special states, the conditions obtained in terms of conditional entropies can simplify the characterization.
VII.2.1 Weyl states in two qubits
Here, we investigate the faithful entanglement of the two-qubit Weyl state. Using Theorem(4), and the result from [54], we obtain the following results for the two-qubit Weyl state.
Observation 4.
If the conditional Rényi $2-$entropy (CR2222E) of two-qubit Weyl state (i.e., $\rho^{weyl}_{2}$) is negative, then $\rho^{weyl}_{2}$ is faithful entangled state.
Using Theorem(5) and the result from [54], we obtain the following results for the two-qubit Weyl state.
Observation 5.
If two-qubit Weyl state is faithful entangled then the conditional Rényi $2-$entropy (CR2222E) is bounded above by $\log_{2}[\frac{1}{1-R}]$ where $R=|t_{1}||t_{2}|+|t_{1}||t_{3}|+|t_{2}||t_{3}|$ and $10<R<$ 1.
VII.2.2 Isotropic state in two qudits
Here, we investigate the faithful entanglement of two-qudit isotropic states. Using Corollary(6.1) and result from [54], we obtain the following result.
Observation 6.
If conditional von Neumann entropy (CVNE) of two-qudit isotropic state is negative then it is faithful entangled.
From Theorem(7), we have
By doing some algebraic calculations, we obtain that $\frac{1+\sqrt{1+d(d^{2}-d-1)}}{d^{2}}>\frac{1}{d}$ for all $d>1$, and hence $F>\frac{1}{d}$. So, we conclude the following
$\text{CR$2$E}(\rho_{d}^{iso})<0\Rightarrow F>\frac{1}{d}.$ (36)
Now, using Equation(36) and the result from [54], we obtain the following result for two-qudit isotropic state.
Observation 7.
If conditional Renyi 2-entropy (CR2222E) of isotropic state with fully entangled fraction $F$ in $d\otimes d$ system is negative then the state is faithful.
VII.2.3 Generalized Bell Diagonal States
Here, we investigate the faithful entanglement for the generalized Bell diagonal state $\rho$. Using Theorem(13) and the result from [54], we see the relation between faithful entanglement and von Neumann entropy and obtain the following results for generalized Bell diagonal state.
Observation 8.
If the generalized Bell diagonal state $\rho$ is faithful then $S(A|B)<Y$ where $Y=\log_{2}\left(\frac{d^{F-1}}{\prod_{i=0}^{d^{2}-2}p_{i}^{p_{i}}}\right),p_{i}>0$.
Now, using Theorem(15) and the result from [54], we see the relation between faithful entanglement and the conditional Rényi $\alpha$- entropy (CRAE), and obtain the following results for generalized Bell diagonal state.
Observation 9.
If the generalized Bell diagonal state $\rho$ is faithful then $S_{\alpha}(A|B)<\frac{1}{1-\alpha}\log_{2}\left(\frac{1+d^{\alpha}Z}{d}\right)$ where $Z=\sum_{i=0}^{d^{2}-2}p_{i}^{\alpha}$ i.e., sum of all $p_{i}^{\alpha}$ except the term for largest $p_{i}$.
VII.3 Entropic uncertainty relation
Consider two incompatible measurements $X$ and $Y$ such that $H(X)$ and $H(Y)$ are Shannon entropies corresponding to the outcomes of measurements when $X$ and $Y$ are performed on a quantum state. Then the entropic uncertainty relation [55] is given by [56, 57]
$H(X)+H(Y)\geq\log_{2}\frac{1}{c},$ (37)
where $c=\text{max}_{i,j}|\langle x_{i}|y_{j}\rangle|^{2}$ such that $|x_{i}\rangle$ and $|y_{j}\rangle$ are eigenvector of $X$ and $Y$, respectively.
The operational interpretation of the above Equation(37) can be understood through the so-called guessing game [58]. However, the uncertainty relation above is no longer valid when one of the parties has access to quantum memory. The entropic uncertainty relation in the quantum memory scenario is stated as [58]
$H(X|B)+H(Y|B)\geq\log_{2}\frac{1}{c}+S(A|B),$ (38)
where $H(X|B)$ or $H(Y|B)$ denote the amount of uncertainty of measurement $X$ or $Y$ given the information stored in quantum memory $B$. Here, from Equation(38), we observe that if $S(A|B)$ (i.e., CVNE) is negative, then the state prepared by Bob overcomes the original uncertainty bound of Equation(37).
Now, we investigate the entropic uncertainty relation for the two-qudit Werner state. By using Theorem(8) and the Equation(38), we obtain the following observations.
Observation 10.
If $d$ is even or odd and $\frac{1}{d}\leq x<1$, then
$H(X|B)+H(Y|B)\geq\log_{2}\frac{1}{c}+\frac{1}{2}\log_{2}(\frac{\Gamma}{F^{\gamma}})$
where $\gamma=1+x$ and $\Gamma=\left(\frac{d^{2}-d}{1-x}\right)^{1-x}\frac{1}{d^{2}}$.
Now, if we consider $F\leq\frac{1}{d}$ in Observation(10), then we obtain the following lower bounds
$H(X|B)+H(Y|B)\geq\log_{2}\frac{1}{c}+\frac{1}{2}\log_{2}(\Gamma d^{\gamma})$
where $\gamma=1+x$ and $\Gamma=\left(\frac{d^{2}-d}{1-x}\right)^{1-x}\frac{1}{d^{2}}$. In a similar way, by using Theorem(9) and the Equation(38), we obtain the following observation for the two-qudit Werner state.
Observation 11.
If $d$ is even and $-1<x<\frac{1}{d}$, then
$H(X|B)+H(Y|B)\geq\log_{2}\frac{1}{c}+\frac{1}{2}\log_{2}(\frac{\Gamma}{F^{\gamma}})$
where $\gamma=1-x$ and $\Gamma=\left(\frac{d^{2}+d}{1+x}\right)^{1+x}\frac{1}{d^{2}}$.
In Observation(11), if we consider $F\leq\frac{1}{d}$, then we obtain the following lower bounds
$H(X|B)+H(Y|B)\geq\log_{2}\frac{1}{c}+\frac{1}{2}\log_{2}(\Gamma d^{\gamma})$
where $\gamma=1-x$ and $\Gamma=\left(\frac{d^{2}+d}{1+x}\right)^{1+x}\frac{1}{d^{2}}$.
In both observations we are also able to find the lower bounds of the entropic uncertainty relation for the states in the class of two-qudit Werner states that are not useful for teleportation.